OpenStudy (andriod09):

Guys, please help. I suck at these problems, on my finals' review, question is in the comments.

OpenStudy (anonymous):

ok I will help but you must help me

OpenStudy (andriod09):

My question: \[\text {One number is 3 more than another. Twice the smaller is equal to 5 times the larger}\] \[\text {decreased by 3. Find the numbers.}\]

OpenStudy (anonymous):

ok i am on it

OpenStudy (anonymous):

Part 4: Translate and solve: 25. 6 less than twice a number is ten Answer : 2n - 6 = 10 (answer: n = 8) 26. The quotient of a number and 4 plus 3 is 8 Answer: n/4 + 3 = 8 (answer: n = 29) 27. 4 less than three times a number is -28 Answer: 3n - 4 = -28 (answer: n = -8) 28. ...

OpenStudy (anonymous):

??? First thing to do is represent, it's always the first thing to do in word problems ^.^

OpenStudy (andriod09):

@JordanFeinstein Please don't jsut copy/paste the answers, as it's concidered cheating according to the CoC, which can be found at: http://www.openstudy.com/code-of-conduct @PeterPan That's the problem, I suck at setting these problems up.

OpenStudy (anonymous):

okay, what's your favourite letter? LOL I'm just kidding, I don't care what your favourite letter is, I'll use x anyway :P Anyway seriously... take things one at a time... -One number...- (let's call this x) -...is three more than than another number- everything ok (so far) ? ^.^

OpenStudy (andriod09):

So then it would be: \[x + (3 \times x)\] then what would be next?

OpenStudy (anonymous):

"three MORE THAN" should be a cue for addition and not multiplication ;)

OpenStudy (anonymous):

question: are we supposed to be working with two variables here or just one?

OpenStudy (andriod09):

Just one variable.

OpenStudy (anonymous):

okay :) So you have two missing numbers, one of them is x, the other one, x is bigger than that other one, by 3, aye? :D

OpenStudy (dumbcow):

its asking for 2 numbers, a smaller and a larger 2 variables :)

OpenStudy (andriod09):

this is the problem: One number is 3 more than another. Twice the smaller is equal to 5 times the larger decreased by 3. Find the numbers. right now, we have \[x + (3+x)\] right?

OpenStudy (anonymous):

Okay, you know what, let's let x be the smaller number instead...

OpenStudy (anonymous):

So if x is the smaller number, and the bigger number is bigger than the smaller number BY THREE, then the bigger number should be x + 3, right? :)

OpenStudy (andriod09):

Aye, but its the same thing just put a different way.

OpenStudy (anonymous):

Yup :) So let's get back to it... -twice the smaller number- How do you represent that? Remember, x is the smaller number... ^.^

OpenStudy (andriod09):

I have no clue. >.<

OpenStudy (anonymous):

Twice is multiplying by 2 :) Twice the smaller number --> 2x okay? ^.^

OpenStudy (andriod09):

Aye, so then you would have: \[2x+(3+x)\] Is that right?

OpenStudy (anonymous):

Not quite :) Let's not get ahead of ourselves yet :D

OpenStudy (anonymous):

-twice the smaller number- (that's 2x) -is equal to- (2x = ...) -five times the larger number- Okay, so how to represent 5 times the larger number? (remember, the larger number is x+3)

OpenStudy (andriod09):

erm. Is it: \[2x=5\times (x+3)\] by any chance?

OpenStudy (anonymous):

yes, but ... at the end, we're not yet finished... -twice the smaller number is equal to five times the larger number DECREASED BY THREE- So it should be \[2x = 5\times (x+3) \color{red}{-3}\] Because of the decreased by three bit at the end...

OpenStudy (anonymous):

And here, you're ready to solve for x ^.^

OpenStudy (anonymous):

Because we have successfully represented the entire problem--

OpenStudy (anonymous):

everything all right? ^.^

OpenStudy (andriod09):

Where'd the little -3 come from? Sorry, didn't go during the lag.

OpenStudy (anonymous):

the little -3 comes from the bit at the end that says DECREASED BY THREE ^.^

OpenStudy (anonymous):

Decreased means subtract ^.^

OpenStudy (andriod09):

*grumbles to him self lowly thinking of how mad he is because he didn't see that* Well then. Lets get to it, shall we?

OpenStudy (anonymous):

Lead the way :D

OpenStudy (anonymous):

Why don't you start with distributing this part \[2x = \color{blue}{5\times(x+3)}-3\]

OpenStudy (andriod09):

Err. I don't know what i'm doing. And I have to get to my ASL class. dang it, TTYL talk to you tomorrow.

OpenStudy (anonymous):

aww, and we were just getting to the fun part... oh well :)

OpenStudy (andriod09):

Tell me about it. >.<

OpenStudy (anonymous):

Have fun ^.^